OpenStudy (anonymous):

The perimeter of a triangle is 60 inches. If the sides of the triangle are 2x, 3x, and 5x, find the length of the shortest side. A.) 1 ft B.) 2 in. C.) 4 in. D.) 6 in.

Vocaloid (vocaloid):

perimeter = 60, which is the sum of all the sides therefore 2x + 3x + 5x = 60 solve for x, then find 2*x

OpenStudy (anonymous):

Can you help me out with that?

OpenStudy (anonymous):

The shortest side length is 12.

Vocaloid (vocaloid):

2x + 3x + 5x = 60 combine like terms and solve for x

OpenStudy (anonymous):

I got 6. is that correct?

Vocaloid (vocaloid):

yes, x = 6, so what is 2*x?

OpenStudy (anonymous):

{2x,3x,5x}/.x->6 , {12,18,30}

OpenStudy (anonymous):

But there is no answer 12!?

OpenStudy (anonymous):

did you solve it?

OpenStudy (anonymous):

all you do is solve for x

OpenStudy (anonymous):

I solved for x. and I got 12. but there isn't an answer 12!

OpenStudy (anonymous):

no add all the sides together

Vocaloid (vocaloid):

hm, that's odd, the answer should be 12 but it's not one of the answer choices in that case, I believe the problem has a typo

Vocaloid (vocaloid):

I would put D for now (even though it's not right) but be sure to tell your teacher about it

OpenStudy (anonymous):

no you are going at it wrong @Vocaloid

Vocaloid (vocaloid):

how am I wrong? 2x + 3x + 5x = 60 x = 6 shortest side = 2x = 2*6 = 12...

OpenStudy (anonymous):

woah nevermind. my bad. wtf man. yea its your question @cheyennerae1605